STRUCTURAL STUDY OF POLAR LIQUIDS THROUGH DIELECTRIC STUDY USING TIME DOMAIN REFLECTOMETERY

\underline{SURESH C.MEHROTRA}, Department of Computer Science Information Technology,Dr. Babasaheb Ambedkar Marathwada University, Aurangabad 431 004.India.

The microwave group at Dr. Babasaheb Ambedkar Marathwada University has reported dielectric relaxation studies on many binary polar liquids The results have been explained qualitatively by assuming formation of multimers in the binary polar liquid. However, no attempt has been yet made to explain the result quantitatively by any group. There is usual practice in the field to explain the experimental dielectric results by assuming their breaking of multimers or formation of multimers. The objective of the present talk is to provide information about experimental aspects of Time Domain Methods along with theoretical models to extract information regarding liquid structure.